[ Ben asks Amy how old she is and she says: "if you subrtact 30 years from the triple of years I had last year, you will get half of my age now", how old is Amy? ]
Answer by TimothyLamb(4379)

3(a - 1) - 30 = (1/2)a
3a - 3 - 30 = (1/2)a
3a - 33 = (1/2)a
(6/2)a - (1/2)a = 33
(5/2)a = 33
a = (2/5)33
---
answer:
Amy is 13.2 years old
that's 13 years, 2 months, and 12 days old
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php